OpenStudy (annie96):

1. Solve. 9x = 3x – 36 (Points : 1) –6 –4 4 6 Question 2.2. Solve. 1/2(21-3a)=6-a (Points : 1) all real numbers no solution –1 0 Question 3.3. Solve. 4(a + 3) = 12 + 4a (Points : 1) –3 –1 no solution all real numbers

OpenStudy (a1234):

Can you give your answers? I'll check them.

OpenStudy (anonymous):

9x = 3x – 36 9x-3x = -36 6x = -36 x = -6

OpenStudy (anonymous):

I'll show you how to work them.

OpenStudy (annie96):

1.c 2.b 3.b

OpenStudy (anonymous):

1/2(21-3a)=6-a 21-3a = 2(6-a) 21-3a = 12 -2a 21-12 = -2a+3a 9 = a

OpenStudy (anonymous):

4(a + 3) = 12 + 4a 4a+12 = 12 + 4a all real numbers

OpenStudy (annie96):

the answer to number 2 is c

OpenStudy (anonymous):

I'm not sure you're right Annie96